# The Landscape of Agentic Reinforcement Learning for LLMs: A Survey

### v1

@@ 4.5 RL in Vision Agents
The success of DeepSeek-R1 [31] has sparked widespread interest in applying RL to incentivize long-form reasoning behavior, encouraging LVLMs to produce extended CoT sequences that improve visual perception and understanding [194]. This research trajectory has evolved from early work that simply adapted R1-style objectives to the vision domain—aimed primarily at enhancing passive perception [206, 207, 208, 200, 201, 358, 359, 360]—toward the now-popular paradigm of active perception, or “thinking with images” [196]. The key transition lies in moving from text-only CoT that references an image once, to interactive, visually grounded reasoning, achieved through (i) grounding [361, 218, 219, 220, 221, 222], (ii) agentic tool use [225, 226, 227, 228, 205, 89], and (iii) visual imagination via sketching or generation [229, 232, 233]. Beyond text-only outputs, many vision tasks—such as scene understanding—require structured predictions like bounding boxes, masks, and segmentation maps. To begin with, Visual-RFT [205] uses IoU with confidence as a verifiable reward for bounding-box outputs, while Vision-R1 [208] incorporates precision and recall as localization rewards. Extending this idea, [362] applies GRPO to segmentation tasks, combining soft and strict rewards with bounding-box IoU and L1 loss, and point-wise L1 distance. VLM-R1 [200] employs mean Average Precision (mAP) as a reward to explicitly incentivize detection and localization capabilities in LVLMs. Finally, R1-SGG [363] introduces three variants of GRPO rewards for scene-graph matching—ranging from hard rewards based on text matching and IoU to softer rewards computed via text-embedding dot products. RL has also been widely applied to image generation, particularly through its integration with diffusion and flow models—for example, RePrompt [364], Diffusion-KTO [365], Flow-GRPO [366], and GoT-R1 [367]. Beyond diffusion-based approaches, RL has been leveraged for autoregressive image generation, where it improves coherence, fidelity, and controllability by directly optimizing task- or user-specific reward signals [368, 233, 369].
@@
Following the same spirit, numerous works have extended GRPO variants to the video domain [370, 371, 372] to enhance temporal reasoning [373, 374, 375, 376, 377]. TW-GRPO [378] introduces a token-weighted GRPO framework that emphasizes high-information tokens to generate more focused reasoning chains and employs soft, multi-choice rewards for lower-variance optimization. EgoVLM [379] combines keyframe-based rewards with direct GRPO training to produce interpretable reasoning traces tailored for egocentric video. DeepVideo-R1 reformulates the GRPO objective as a regression task [373], while VideoChat-R1 demonstrates that reinforcement finetuning (RFT) can be highly data-efficient for task-specific video reasoning improvements [374]. TinyLLaVA-Video-R1 explores scaling RL to smaller video LLMs [380], and [381] introduces infrastructure and a two-stage pipeline (CoT-SFT + RL) to support large-scale RL for long videos. Additional efforts have also extended RL for embodied video reasoning tasks [382]. A similar trend is observed in video generation, where RL is applied to improve temporal coherence, controllability, and semantic alignment. Key examples include DanceGRPO [383], GAPO [384], GRADEO [385], InfLVG [386], Phys-AR [387], VideoReward [388], TeViR [389], and InstructVideo [390].
@@
RL has also been widely adopted to advance 3D understanding [391, 392, 393, 394, 395, 396] and generation [397, 398, 399]. MetaSpatial [400] introduces the first RL-based framework for 3D spatial reasoning, leveraging physics-aware constraints and rendered-image evaluations as rewards during training. Scene-R1 [401] learns to reason about 3D scenes without point-wise 3D supervision, while SpatialReasoner [402] introduces shared 3D representations that unify perception, computation, and reasoning stages. In the domain of 3D generation, RL has been applied to improve text-to-3D alignment and controllability. Notable efforts include DreamCS [403], which aligns generation with human preferences; DreamDPO [404] and DreamReward [405], which optimize 3D generation using 2D reward signals; and Nabla-R2D3 [406], which further refines 3D outputs with reinforcement-driven objectives.
4.6 RL in Embodied Agents
While traditional agents are typically developed for general-purpose vision or language tasks, extending these capabilities to embodied agents requires a comprehensive understanding of real-world visual environments and the capacity to reason across modalities. Such competencies are essential for perceiving complex physical contexts and executing goal-directed actions conditioned on high-level instructions, forming a foundational element of agentic LLMs and MLLMs. In instruction-driven embodied scenarios, RL is often employed as a post-training strategy. A common pipeline begins with a pre-trained vision-language-action (VLA) model [407, 408, 409, 410] obtained through imitation learning under teacher-forcing supervision. This model is then embedded into an interactive agent that engages with the environment to collect reward signals. These rewards guide the iterative refinement of the policy, supporting effective exploration, improving sample efficiency, and enhancing the model’s generalization capabilities across diverse real-world conditions. RL in VLA frameworks [411, 412, 413, 414] can be broadly categorized into two classes: navigation agents, which emphasize spatial reasoning and locomotion in complex environments, and manipulation agents, which focus on the precise control of physical objects under diverse and dynamic constraints.
@@ 4.6 RL in Embodied Agents
For navigation agents, planning is the central capability. Reinforcement learning is employed to enhance the VLA model’s ability to predict and optimize future action sequences. A common strategy [415] is to integrate traditional robotics-style RL, using step-wise directional rewards, directly into VLA-based navigation frameworks. Some approaches operate at the trajectory level. VLN-R1 [413] aligns predicted and ground-truth paths to define trajectory-level rewards, and applies GRPO, following DeepSeek-R1, to improve predictive planning. OctoNav-R1 [416] also leverages GRPO but focuses on reinforcing internal deliberation within the VLA model, promoting a thinking-before-acting paradigm that enables more anticipatory and robust navigation. S2E [417] introduces a reinforcement learning framework that augments navigation foundation models with interactivity and safety, combining video pretraining with RL to achieve superior generalization and performance on the NavBench-GS benchmark.
@@
Manipulation agents, typically involving robotic arms, require fine-grained control for executing structured tasks under diverse conditions. In this context, RL is employed to enhance the instruction-following and trajectory prediction capabilities of VLA models, especially to improve generalization across tasks and environments. RLVLA [418] and VLA-RL [412] adopt pre-trained VLMs as evaluators, using their feedback to assign trajectory-level rewards for VLA policy refinement. These methods establish an online RL framework that effectively improves manipulation performance and demonstrates favorable scaling properties. TGRPO further [419] incorporates GRPO into manipulation tasks by defining rule-based reward functions over predicted trajectories. This enables the VLA model to generalize to unseen scenarios and improves its robustness in real-world deployment. VIKI-R [420] complements this with a unified benchmark and two-stage framework for multi-agent embodied cooperation, combining Chain-of-Thought fine-tuning with multi-level RL to enable compositional coordination across diverse embodiments.
@@
4.7 RL in Multi-Agent Systems
Large Language Model (LLM)-based Multi-agent Systems (MAS) comprise multiple autonomous agents collaborating to solve complex tasks through structured interaction, coordination, and memory management. Early static and hand-designed MAS such as CAMEL and MetaGPT [421, 268] explored role specialization and task decomposition, while debate-based frameworks such as MAD and MoA [422, 423] enhanced reasoning via collaborative refinement. Subsequent multi-agent research has shifted to proposing optimizable cooperative systems, which enable MAS to not only dynamically adjust coordination patterns but also directly enhance agent-level reasoning and decision-making strategies. Table 8 summarizes the main body of works discussed in this section.
@@ 4.7 RL in Multi-Agent Systems
In the RL-free self-evolving setting, foundation models cannot be directly optimized; instead, system evolution is driven by mechanisms such as symbolic learning [424], dynamic graph optimization [425, 426, 427], and workflow rewriting [428, 429, 430]. These methods improve the coordination and adaptability within MAS, but cannot directly update the parameters of foundation models. MALT [154] employs a heterogeneous multi-agent search tree to generate large-scale labeled trajectories, fine-tuning agents via a combination of Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) from both successful and failed reasoning paths.
@@
[…]
@@
[…]
MALT [154] DPO -
[…]
ACC-Collab [146] DPO -
[…]
MLPO [435] MLPO -
[…]
MMedAgent-RL [438] GRPO -
[…]
RLCCF [440] GRPO -
MAGRPO [441] MAGRPO -
RL-Based Multi-Agent Training
@@
MARFT [433] formalizes a reinforcement fine-tuning framework for MAS with mathematical guarantees and empirical validation. MAGRPO [441] formalizes multi-LLM cooperation as a Dec-POMDP problem and introduces a multi-agent variant of GRPO, which enables joint training of LLM agents in MAS while maintaining decentralized execution. MAPoRL [434] extends MAD by verifying debate responses and using validation outcomes as RL rewards to improve collaborative reasoning. RLCCF [440] is a self-supervised multi-agent RL framework that leverages self-consistency-weighted ensemble voting to generate pseudo-labels and collaboratively optimize individual model policies via GRPO, boosting both individual and collective reasoning accuracy. MLPO [435] introduces a hierarchical paradigm in which a central LLM leader learns, via RL, to synthesize and evaluate peer agent outputs without auxiliary value networks. ReMA [436] separates reasoning into a meta-thinking agent and an execution agent, jointly trained under aligned RL objectives with parameter sharing. FlowReasoner [437] designs query-level meta-agents optimized through RL with multi-dimensional rewards (performance, complexity, efficiency) guided by execution feedback. LERO [442] combines MARL with LLM-generated hybrid rewards and evolutionary search to improve credit assignment and partial observability handling in cooperative tasks. CURE [275] focuses on code generation, jointly training a code generator and unit tester via RL to produce richer reward signals, achieving strong generalization across diverse coding benchmarks. MMedAgent-RL [438] introduces a reinforcement learning-based multi-agent framework for medical VQA, where dynamically coordinated general practitioners and specialists collaboratively reason with curriculum-guided learning, significantly outperforming existing Med-LVLMs and achieving more human-like diagnostic behavior. Chain-of-Agents (COA) [439] is an end-to-end paradigm where a single LLM simulates multi-agent collaboration by dynamically orchestrating role-playing and tool-using agents; this is achieved through multi-agent distillation (converting trajectories from state-of-the-art multi-agent systems into training data) and agentic reinforcement learning with carefully designed reward functions, resulting in Agent Foundation Models (AFMs). SPIRAL [443] presents a fully online, multi-turn, multi-agent self-play reinforcement learning framework for LLMs in zero-sum games, employing a shared policy with role-conditioned advantage estimation (RAE) to stabilize learning, and demonstrates that gameplay fosters transferable reasoning skills that significantly improve mathematical and general reasoning benchmarks.

### v4

@@ 4.5 Vision Agents
The success of DeepSeek-R1 (DeepSeek-AI et al., 2025) has sparked widespread interest in applying RL to incentivize long-form reasoning behavior, encouraging LVLMs to produce extended CoT sequences that improve visual perception and understanding (Shao et al., 2024a). This research trajectory has evolved from early work that simply adapted R1-style objectives to the vision domain—aimed primarily at enhancing passive perception (Tan et al., 2025a; Li et al., 2025t; Huang et al., 2025c; Shen et al., 2025a; Peng et al., 2025; Xia et al., 2025a; Yang et al., 2025c; Gao et al., 2025a)—toward the now-popular paradigm of active perception, or “thinking with images” (Su et al., 2025c). The key transition lies in moving from text-only CoT that references an image once, to interactive, visually grounded reasoning, achieved through (i) grounding (Li et al., 2025e; Nagaraja et al., 2016; Mao et al., 2016; Fan et al., 2025c; Chung et al., 2025; Cao et al., 2025a), (ii) agentic tool use (Zhao et al., 2025d; Huang et al., 2025d; Wu et al., 2025g; Su et al., 2025b; Liu et al., 2025v; Su et al., 2025a), and (iii) visual imagination via sketching or generation (Xu et al., 2025e; Duan et al., 2025; Jiang et al., 2025b). Beyond text-only outputs, many vision tasks—such as scene understanding—require structured predictions like bounding boxes, masks, and segmentation maps. To begin with, Visual-RFT (Liu et al., 2025v) uses IoU with confidence as a verifiable reward for bounding-box outputs, while Vision-R1 (Huang et al., 2025c) incorporates precision and recall as localization rewards. Extending this idea, Liu et al. (2025p) applies GRPO to segmentation tasks, combining soft and strict rewards with bounding-box IoU and L1 loss, and point-wise L1 distance. VLM-R1 (Shen et al., 2025a) employs mean Average Precision (mAP) as a reward to explicitly incentivize detection and localization capabilities in LVLMs. Finally, R1-SGG (Chen et al., 2025m) introduces three variants of GRPO rewards for scene-graph matching—ranging from hard rewards based on text matching and IoU to softer rewards computed via text-embedding dot products. RL has also been widely applied to image generation, particularly through its integration with diffusion and flow models—for example, RePrompt (Wu et al., 2025f), Diffusion-KTO (Li et al., 2024c), Flow-GRPO (Liu et al., 2025d), and GoT-R1 (Duan et al., 2025). Beyond diffusion-based approaches, RL has been leveraged for autoregressive image generation, where it improves coherence, fidelity, and controllability by directly optimizing task- or user-specific reward signals (Wang et al., 2025f; Jiang et al., 2025b; Yuan et al., 2025a).
@@
Following the same spirit, numerous works have extended GRPO variants to the video domain (Cheng et al., 2024; Feng et al., 2024b; Maaz et al., 2023) to enhance temporal reasoning (Park et al., 2025b; Li et al., 2025k; Zhu et al., 2025c; Liao et al., 2025c; Ouyang, 2025). TW-GRPO (Dang et al., 2025) introduces a token-weighted GRPO framework that emphasizes high-information tokens to generate more focused reasoning chains and employs soft, multi-choice rewards for lower-variance optimization. EgoVLM (Vinod et al., 2025) combines keyframe-based rewards with direct GRPO training to produce interpretable reasoning traces tailored for egocentric video. DeepVideo-R1 reformulates the GRPO objective as a regression task (Park et al., 2025b), while VideoChat-R1 demonstrates that reinforcement fine-tuning (RFT) can be highly data-efficient for task-specific video reasoning improvements (Li et al., 2025k). TinyLLaVA-Video-R1 explores scaling RL to smaller video LLMs (Zhang et al., 2025t), and (Chen et al., 2025j) introduces infrastructure and a two-stage pipeline (CoT-SFT + RL) to support large-scale RL for long videos. Additional efforts have also extended RL for embodied video reasoning tasks (Zhao et al., 2025b). A similar trend is observed in video generation, where RL is applied to improve temporal coherence, controllability, and semantic alignment. Key examples include DanceGRPO (Xue et al., 2025), GAPO (Zhu et al., 2025a), GRADEO (Mou et al., 2025), InfLVG (Fang et al., 2025b), Phys-AR (Lin et al., 2025c), VideoReward (Liu et al., 2025e), TeViR (Chen et al., 2025i), and InstructVideo (Yuan et al., 2024).
@@
RL has also been widely adopted to advance 3D understanding (Hong et al., 2023; Xu et al., 2024b; Deng et al., 2024a; Chen et al., 2024a; Zhou et al., 2023; Chen et al., 2024d) and generation (Wang et al., 2024i; Yin et al., 2025; Siddiqui et al., 2024). MetaSpatial (Pan & Liu, 2025) introduces the first RL-based framework for 3D spatial reasoning, leveraging physics-aware constraints and rendered-image evaluations as rewards during training. Scene-R1 (Yuan et al., 2025d) learns to reason about 3D scenes without point-wise 3D supervision, while SpatialReasoner (Ma et al., 2025c) introduces shared 3D representations that unify perception, computation, and reasoning stages. In the domain of 3D generation, RL has been applied to improve text-to-3D alignment and controllability. Notable efforts include DreamCS (Zou et al., 2025), which aligns generation with human preferences; DreamDPO (Zhou et al., 2025f) and DreamReward (Ye et al., 2024), which optimize 3D generation using 2D reward signals; and Nabla-R2D3 (Liu et al., 2025i), which further refines 3D outputs with reinforcement-driven objectives.
4.6 Embodied Agents
Embodied agents encompass a broad family of systems that perceive a structured environment and act within it, ranging from vision-language-action (VLA) models to language-driven open-ended agents. While many recent systems focus on VLA settings that require grounding in real-world visual observations, all embodied agents must integrate perception, reasoning, and action to operate effectively in complex physical or simulated environments and to execute goal-directed behaviors conditioned on high-level instructions. These competencies form a foundational component of agentic LLMs and MLLMs in embodied scenarios. In instruction-driven embodied scenarios, RL is often employed as a post-training strategy. A common pipeline begins with a pre-trained vision-language-action (VLA) model (Kim et al., 2024; Black et al., 2024; Team et al., 2025a; Liao et al., 2025b) obtained through imitation learning under teacher forcing supervision. This model is then embedded into an interactive agent that engages with the environment to collect reward signals. These rewards guide the iterative refinement of the policy, supporting effective exploration, improving sample efficiency, and enhancing the model’s generalization capabilities across diverse real-world conditions. RL in VLA frameworks (SimpleVLA-RL Team, 2025; Lu et al., 2025a; Qi et al., 2025; Song et al., 2025e) can be broadly categorized into two classes: navigation agents, which emphasize spatial reasoning and locomotion in complex environments, and manipulation agents, which focus on the precise control of physical objects under diverse and dynamic constraints.
@@ 4.6 Embodied Agents
For navigation agents, planning is the central capability. Reinforcement learning is employed to enhance the VLA model’s ability to predict and optimize future action sequences. A common strategy (Zhao et al., 2025c) is to integrate traditional robotics-style RL, using step-wise directional rewards, directly into VLA-based navigation frameworks. Some approaches operate at the trajectory level. VLN-R1 (Qi et al., 2025) aligns predicted and ground-truth paths to define trajectory-level rewards, and applies GRPO, following DeepSeek-R1, to improve predictive planning. OctoNav-R1 (Gao et al., 2025a) also leverages GRPO but focuses on reinforcing internal deliberation within the VLA model, promoting a thinking-before-acting paradigm that enables more anticipatory and robust navigation. S2E (He et al., 2025) introduces a reinforcement learning framework that augments navigation foundation models with interactivity and safety, combining video pretraining with RL to achieve superior generalization and performance on the NavBench-GS benchmark.
@@
Manipulation agents, typically involving robotic arms, require fine-grained control for executing structured tasks under diverse conditions. In this context, RL is employed to enhance the instruction-following and trajectory prediction capabilities of VLA models, especially to improve generalization across tasks and environments. RLVLA (Liu et al., 2025f) and VLA-RL (Lu et al., 2025a) adopt pre-trained VLMs as evaluators, using their feedback to assign trajectory-level rewards for VLA policy refinement. These methods establish an online RL framework that effectively improves manipulation performance and demonstrates favorable scaling properties. TGRPO  (Chen et al., 2025k) further incorporates GRPO into manipulation tasks by defining rule-based reward functions over predicted trajectories. This enables the VLA model to generalize to unseen scenarios and improves its robustness in real-world deployment. VIKI-R (Kang et al., 2025a) complements this with a unified benchmark and two-stage framework for multi-agent embodied cooperation, combining Chain-of-Thought fine-tuning with multi-level RL to enable compositional coordination across diverse embodiments.
@@
Case Study: Voyager.
Beyond these general challenges in embodied RL, Voyager (Wang et al., 2024a), a language-driven open-ended embodied agent, illustrates how planning, skill acquisition, and RL-based curriculum learning can be integrated in practice. The agent explores Minecraft using an iterative loop: it generates a plan, interacts with the environment, extracts reusable skills from successful trajectories, and stores them in a growing skill library. A curriculum scheduler selects new tasks based on the agent’s current skill set, while RL objectives guide which behavior should be committed as skills and when to refine or discard them. This creates a self-improving cycle in which planning, environmental interaction, memory, and RL-driven curriculum optimization are tightly coupled.
4.7 Multi-Agent Systems
Large Language Model (LLM)-based Multi-agent Systems (MAS) comprise multiple autonomous agents collaborating to solve complex tasks through structured interaction, coordination, and memory management. Early static and hand-designed MAS such as CAMEL and MetaGPT (Li et al., 2023a; Hong et al., 2024b) explored role specialization and task decomposition, while debate-based frameworks such as MAD and MoA (Wang et al., 2025g; Liang et al., 2024) enhanced reasoning via collaborative refinement. Subsequent multi-agent research has shifted to proposing optimizable cooperative systems, which enable MAS to not only dynamically adjust coordination patterns but also directly enhance agent-level reasoning and decision-making strategies. Table 8 summarizes the main body of works discussed in this section.
@@ 4.7 Multi-Agent Systems
In the RL-free self-evolving setting, foundation models cannot be directly optimized; instead, system evolution is driven by mechanisms such as symbolic learning (Zhou et al., 2024c), dynamic graph optimization (Zhuge et al., 2024; Ma et al., 2025d; Zhou et al., 2025b), and workflow rewriting (Hu et al., 2025d; Zhang et al., 2025i; h). These methods improve the coordination and adaptability within MAS, but cannot directly update the parameters of foundation models.
@@
CAMEL (Li et al., 2023a) - GitHub [Uncaptioned image]HuggingFace
MetaGPT (Hong et al., 2024b) - GitHub
MAD (Liang et al., 2024) - GitHub
MoA (Wang et al., 2025g) - GitHub
AFlow (Zhang et al., 2025i) - GitHub
@@
GPTSwarm (Zhuge et al., 2024) policy gradient GitHub Website
MaAS (Zhang et al., 2025f) policy gradient GitHub
G-Designer (Zhang et al., 2025g) policy gradient GitHub
Optima (Chen et al., 2025e) DPO GitHub
DITS (Shi et al., 2025a) DPO -
MALT (Motwani et al., 2025) DPO -
MARFT (Liao et al., 2025a) MARFT GitHub
ACC-Collab (Estornell et al., 2025b) DPO -
MAPoRL (Park et al., 2025a) PPO GitHub
MLPO (Estornell et al., 2025a) MLPO -
ReMA (Wan et al., 2025b) MAMRP GitHub
FlowReasoner (Gao et al., 2025b) GRPO GitHub
CURE (Wang et al., 2025q) rule-based RL GitHub [Uncaptioned image]HuggingFace
MMedAgent-RL (Xia et al., 2025b) GRPO -
Chain-of-Agents (Li et al., 2025h) DAPO GitHub [Uncaptioned image]HuggingFace
RLCCF (Yuan et al., 2025b) GRPO -
MAGRPO (Liu et al., 2025l) MAGRPO -
4.7.1 RL-Driven Optimization of Non-Parametric Coordination Modules
These approaches keep agent parameters frozen while using RL to optimize external coordination structures such as communication topologies, routing policies, or workflow graphs. Methods such as GPTSwarm, MaAS, and G-Designer (Zhuge et al., 2024; Zhang et al., 2025f; g) treat MAS coordination as a graph-level policy updated via policy gradient. Because no agent-level gradients exist, credit assignment must operate at the topology or message-routing level. Rewards are typically delayed and sparse—e.g., only final task accuracy—requiring global-to-local credit decomposition or structural priors to avoid collapse.
A key comparison emerges between fixed communication protocols (pre-specified message formats) and learnable protocols. Fixed protocols excel in low-data or highly specialized domains where stability is critical, whereas learnable protocols allow RL to discover efficient emergent communication but require substantially higher sample complexity and careful regularization to prevent overfitting or degenerate conventions.
4.7.2 RL-Driven Optimization of Selected Agent Policies
A second class of systems updates only a subset of agents—often a leader, coordinator, or specialized expert—while keeping others frozen for stability. Representative examples include Optima, DITS, MALT, ACC-Collab (Chen et al., 2025e; Shi et al., 2025a; Motwani et al., 2025; Estornell et al., 2025b). These approaches balance flexibility and scalability: training only a few agents reduces sample complexity and avoids the instability of fully-decoupled credit assignment. MALT (Motwani et al., 2025) employs a heterogeneous multi-agent search tree to generate large-scale labeled trajectories, fine-tuning agents via a combination of Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) from both successful and failed reasoning paths.
Credit assignment in this regime is fundamentally semi-local: rewards emerge from a collective trajectory, but gradients apply only to the optimized agent(s). This requires mechanisms such as role-conditioned DPO (Motwani et al., 2025), local advantage estimation, or counterfactual baselines to prevent reward hijacking by non-updated agents. Empirically, such partial optimization yields better sample efficiency than fully joint multi-agent training while still enabling the emergence of specialized roles.
4.7.3 End-to-End Multi-Agent Reinforcement Learning
@@
Full multi-agent RL jointly trains all agents under a shared or decentralized objective, typically formalized as a Dec-POMDP. Methods such as MAGRPO, MAPoRL, MLPO, ReMA, FlowReasoner, Chain-of-Agents, and SPIRAL (Liu et al., 2025l; Park et al., 2025a; Estornell et al., 2025a; Wan et al., 2025b; Gao et al., 2025b; Li et al., 2025h; Liu et al., 2025b) jointly optimize collaboration and reasoning behaviors, enabling emergent division of labor and communication conventions. For example, MAGRPO (Liu et al., 2025l) formalizes multi‑LLM cooperation as a Dec-POMDP problem and introduces a multi‑agent variant of GRPO, which enables joint training of LLM agents in MAS while maintaining decentralized execution. MAPoRL (Park et al., 2025a) extends MAD by verifying debate responses and using validation outcomes as RL rewards to improve collaborative reasoning. RLCCF (Yuan et al., 2025b) is a self-supervised multi-agent RL framework that leverages self-consistency-weighted ensemble voting to generate pseudo-labels and collaboratively optimize individual model policies via GRPO, boosting both individual and collective reasoning accuracy. ReMA (Wan et al., 2025b) separates reasoning into a meta-thinking agent and an execution agent, jointly trained under aligned RL objectives with parameter sharing. LERO (Wei et al., 2025d) combines MARL with LLM-generated hybrid rewards and evolutionary search to improve credit assignment and partial observability handling in cooperative tasks. CURE (Wang et al., 2025q) focuses on code generation, jointly training a code generator and unit tester via RL to produce richer reward signals, achieving strong generalization across diverse coding benchmarks. MMedAgent-RL (Xia et al., 2025b) introduces a reinforcement learning-based multi-agent framework for medical VQA, where dynamically coordinated general practitioners and specialists collaboratively reason with curriculum-guided learning, significantly outperforming existing Med-LVLMs and achieving more human-like diagnostic behavior. Chain-of-Agents (COA) (Li et al., 2025h) is an end-to-end paradigm where a single LLM simulates multi-agent collaboration by dynamically orchestrating role-playing and tool-using agents; this is achieved through multi-agent distillation (converting trajectories from state-of-the-art multi-agent systems into training data) and agentic reinforcement learning with carefully designed reward functions, resulting in Agent Foundation Models (AFMs). SPIRAL (Liu et al., 2025b) presents a fully online, multi-turn, multi-agent self-play reinforcement learning framework for LLMs in zero-sum games, employing a shared policy with role-conditioned advantage estimation (RAE) to stabilize learning, and demonstrates that gameplay fosters transferable reasoning skills that significantly improve mathematical and general reasoning benchmarks.
However, end-to-end multi-LLM training exacerbates the temporal and structural credit assignment problem because rewards may depend on long multi-turn interaction chains. Solutions include role-conditioned advantage estimation (RAE), hierarchical controller–worker architectures (MLPO, ReMA), and self-play curricula (SPIRAL) that densify reward signals by constructing increasingly challenging interactions. These hierarchical patterns mirror enterprise deployments where a supervisory agent coordinates multiple workers; RL proves particularly effective at learning stable delegation and arbitration strategies under sparse reward settings. Despite their expressiveness, joint MARL approaches face scalability limits: sample complexity grows roughly linearly with the number of agents and quadratically with interaction depth. Algorithms such as MAGRPO and PPO-based MAPoRL mitigate this using centralized critics or value-shared baselines, but achieving scalable credit decomposition remains a central open challenge.